(10pts) Jim Clark borrowed $20,000 from a friendly credit union to buy a car at 6% interest per year compounded monthly for 60 months. The monthly payment for Jim is nearly equal to _______________.

$386.66

$399.99

$410.33

$520.43

(10pts) It is expected that an investment of “P” at time “0” to provide an income of $1,000 every year for the next 15 years. If the investor expects to have a rate of return of 7% per year, what is the value of “P”?

$6,266.66

$9.107.91

$7,107.91

$8,107.91
Answer #1
1) Monthly Payment = Loan amount / Present value of annuity of 1
= $ 20,000.00 / 51.72556075
= $ 386.66
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005)^-60)/0.005 i = 6%/12 = 0.005
= 51.72556075 n = 60
2) Value of "P" = Annual income * Present value of annuity of 1
= $ 1,000.00 * 9.107914005
= $ 9,107.91
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.07)^-15)/0.07 i = 7%
= 9.107914005 n = 15